write the slope intercept equation for the line that passes through (6,-9)and has slope -7/6.
Answer by lwsshak3(11628)

write the slope intercept equation for the line that passes through (6,-9)and has slope -7/6.
**
standard form of equation for a straight line: y-=mx+b, m=slope, b=y-intercept
Given equation:
y=-7x/6+b
solving for b using coordinates of given point (6,-9)
-9=-7*6/6+b
b=-9+7=-2
equation:
y=-7x/6-2